Problem 2.187 Samples of a cast aluminum part are classifed on the basis of surface finish (in microinches) and edge finish. The results of 115 parts are summariaed as follows edge finish excellent good excelent 88 finish good 2 10 Let A denote the event that a sample has excellent surface finish, and let 8 denote probabilities. Round your answers to three decimal the event that a sample has excellent edge finish. If a part is selected at random, determine the following (a) PIA) 0808 (b) P(B)- 0.869 e PA) (d) PIAN8) e) PIAUB) PIA UB) Statistical Tables and Charts
Answer #1

(a)

Edge finish excellent Edge finish good Total
Surface finish excellent 88 5 93
Surface finish good 12 10 22
Total 100 15 115

A = Surface finish excellent

B = Edge finish excellent

P(A)= 93/115 = 0.8087

(b)

P(B) = 100/115 = 0.8696

(c)

P(A') = 1 - P(A) = 0.1913

(d)

115 = 0.7652

(e)

P(AUB)- P(A) +PUB) - PAB)

0.8087 0.8696-0.7652-0.9131

(f)

P(A'igcup B)=P(A')+P(B)-P(A'igcap B)

= 0.1913 + 0.8696 - (12/115)

= 1.0609 - 0.1043 =  0.9566